OpenStudy (anonymous):

Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v. <-20, -42> <-20, 0> <0, 18> <-6, 0>

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (anonymous):

I originally put C and got it wrong so no :/

OpenStudy (anonymous):

Did you type your options rightly? or is A supposed to be -20, -24?

jimthompson5910 (jim_thompson5910):

to find -2u you would multiply everything in vector u by -2 so what is the vector -2u equal to?

OpenStudy (anonymous):

No i copied it straight from the test.. and -2(5(-2)) ? irdk

jimthompson5910 (jim_thompson5910):

u = <5, 6> -2*u = -2*<5, 6> -2*u = <-2*5, -2*6> ... notice how I'm multiplying everything by -2 -2u = <-10, -12>

jimthompson5910 (jim_thompson5910):

do the same for 5v then add up -2u and 5v

OpenStudy (anonymous):

Yeah so then 5v would equal v= <-2,-6> 5*v= 5*<-2,-6> 5*v=<5*(-2)5*(-6)> 5*v= <-10,-30>

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

to add vectors, you just add the corresponding components

jimthompson5910 (jim_thompson5910):

example <3,7> + <-2, 6> = <3+(-2), 7+6> = <1, 13>

OpenStudy (anonymous):

so i would add <-10, -12>+<-10,-30> ??

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so how would i do that ? -10+(-10) and -12+(-30) ?

jimthompson5910 (jim_thompson5910):

yes add up the individual components

OpenStudy (anonymous):

<-20,-42>

jimthompson5910 (jim_thompson5910):

correct. That's your final answer

OpenStudy (anonymous):

omg your a life saver.. wanna help me on some more?